2^x+2 - 2^x = 48

Please explain the steps. Thanks.

Found 2 solutions by jim_thompson5910, scott8148:
Answer by jim_thompson5910(35256)
Answer by scott8148(6628)

by rules of exponents, 2^(x+2)=(2^x)(2^2) ... so 4(2^x)-2^x=48 ... 3(2^x)=48 ... 2^x=16 ... so x=4

2*2*2*2=16